% Generated by IEEEtran.bst, version: 1.14 (2015/08/26)
\begin{thebibliography}{10}
\providecommand{\url}[1]{#1}
\csname url@samestyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand{\BIBentrySTDinterwordspacing}{\spaceskip=0pt\relax}
\providecommand{\BIBentryALTinterwordstretchfactor}{4}
\providecommand{\BIBentryALTinterwordspacing}{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus
  \fontdimen4\font\relax}
\providecommand{\BIBforeignlanguage}[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\typeout{** WARNING: IEEEtran.bst: No hyphenation pattern has been}%
\typeout{** loaded for the language `#1'. Using the pattern for}%
\typeout{** the default language instead.}%
\else
\language=\csname l@#1\endcsname
\fi
#2}}
\providecommand{\BIBdecl}{\relax}
\BIBdecl

\bibitem{hardle1991smoothing}
W.~K. H{\"a}rdle \emph{et~al.}, \emph{Smoothing techniques: with implementation
  in S}.\hskip 1em plus 0.5em minus 0.4em\relax Springer Science \& Business
  Media, 1991.

\bibitem{simonoff2012smoothing}
J.~S. Simonoff, \emph{Smoothing methods in statistics}.\hskip 1em plus 0.5em
  minus 0.4em\relax Springer Science \& Business Media, 2012.

\bibitem{roberts1987digital}
R.~A. Roberts and C.~T. Mullis, \emph{Digital signal processing}.\hskip 1em
  plus 0.5em minus 0.4em\relax Addison-Wesley Longman Publishing Co., Inc.,
  1987.

\bibitem{orfanidis1995introduction}
S.~J. Orfanidis, \emph{Introduction to signal processing}.\hskip 1em plus 0.5em
  minus 0.4em\relax Prentice-Hall, Inc., 1995.

\bibitem{fong2002monte}
W.~Fong, S.~J. Godsill, A.~Doucet, and M.~West, ``Monte carlo smoothing with
  application to audio signal enhancement,'' \emph{IEEE transactions on signal
  processing}, vol.~50, no.~2, pp. 438--449, 2002.

\bibitem{schaub2018flow}
M.~T. Schaub and S.~Segarra, ``Flow smoothing and denoising: Graph signal
  processing in the edge-space,'' in \emph{2018 IEEE Global Conference on
  Signal and Information Processing (GlobalSIP)}.\hskip 1em plus 0.5em minus
  0.4em\relax IEEE, 2018, pp. 735--739.

\bibitem{kenny1998data}
D.~Kenny, D.~Kashy, and N.~Bolger, ``Data analysis,'' in \emph{The handbook of
  social psychology: Vols. 1 and 2,}.\hskip 1em plus 0.5em minus 0.4em\relax
  McGraw-Hill New York, 1998, pp. 233--265.

\bibitem{brandt1998data}
S.~Brandt and S.~Brandt, \emph{Data analysis}.\hskip 1em plus 0.5em minus
  0.4em\relax Springer, 1998.

\bibitem{guo2004functional}
W.~Guo, ``Functional data analysis in longitudinal settings using smoothing
  splines,'' \emph{Statistical methods in medical research}, vol.~13, no.~1,
  pp. 49--62, 2004.

\bibitem{deveaux1999applied}
R.~D. Deveaux, ``Applied smoothing techniques for data analysis,'' 1999.

\bibitem{zin2020smoothing}
M.~A.~M. Zin, A.~S. Rambely, N.~M. Ariff, and M.~S. Ariffin, ``Smoothing and
  differentiation of kinematic data using functional data analysis approach: an
  application of automatic and subjective methods,'' \emph{Applied Sciences},
  vol.~10, no.~7, p. 2493, 2020.

\bibitem{gokcesu2021regret}
K.~Gokcesu and H.~Gokcesu, ``Regret analysis of global optimization in
  univariate functions with lipschitz derivatives,'' \emph{arXiv preprint
  arXiv:2108.10859}, 2021.

\bibitem{saul2004overview}
S.~Carliner, ``An overview of online learning (2nd ed.),'' \emph{European
  Business Review}, vol.~16, 01 2004.

\bibitem{gokcesu2020recursive}
K.~Gokcesu and H.~Gokcesu, ``Recursive experts: An efficient optimal mixture of
  learning systems in dynamic environments,'' \emph{arXiv preprint
  arXiv:2009.09249}, 2020.

\bibitem{yaroshinsky2001smooth}
R.~Yaroshinsky and R.~El-Yaniv, \emph{Smooth online learning of expert
  advice}.\hskip 1em plus 0.5em minus 0.4em\relax Citeseer, 2001.

\bibitem{gokcesu2020generalized}
K.~Gokcesu and H.~Gokcesu, ``A generalized online algorithm for translation and
  scale invariant prediction with expert advice,'' \emph{arXiv preprint
  arXiv:2009.04372}, 2020.

\bibitem{servedio2003smooth}
R.~A. Servedio, ``Smooth boosting and learning with malicious noise,''
  \emph{The Journal of Machine Learning Research}, vol.~4, pp. 633--648, 2003.

\bibitem{gokcesu2021optimal}
K.~Gokcesu and H.~Gokcesu, ``Optimal and efficient algorithms for general
  mixable losses against switching oracles,'' \emph{arXiv preprint
  arXiv:2108.06411}, 2021.

\bibitem{hartikainen2010kalman}
J.~Hartikainen and S.~S{\"a}rkk{\"a}, ``Kalman filtering and smoothing
  solutions to temporal gaussian process regression models,'' in \emph{2010
  IEEE international workshop on machine learning for signal processing}.\hskip
  1em plus 0.5em minus 0.4em\relax IEEE, 2010, pp. 379--384.

\bibitem{gokcesu2021optimally}
K.~Gokcesu and H.~Gokcesu, ``Optimally efficient sequential calibration of
  binary classifiers to minimize classification error,'' \emph{arXiv preprint
  arXiv:2108.08780}, 2021.

\bibitem{chandola2009anomaly}
V.~Chandola, A.~Banerjee, and V.~Kumar, ``Anomaly detection: A survey,''
  \emph{ACM computing surveys (CSUR)}, vol.~41, no.~3, pp. 1--58, 2009.

\bibitem{gokcesu2017online}
K.~Gokcesu and S.~S. Kozat, ``Online anomaly detection with minimax optimal
  density estimation in nonstationary environments,'' \emph{IEEE Transactions
  on Signal Processing}, vol.~66, no.~5, pp. 1213--1227, 2017.

\bibitem{tsopelakos2019sequential}
A.~Tsopelakos, G.~Fellouris, and V.~V. Veeravalli, ``Sequential anomaly
  detection with observation control,'' in \emph{2019 IEEE International
  Symposium on Information Theory (ISIT)}.\hskip 1em plus 0.5em minus
  0.4em\relax IEEE, 2019, pp. 2389--2393.

\bibitem{gokcesu2018sequential}
K.~Gokcesu, M.~M. Neyshabouri, H.~Gokcesu, and S.~S. Kozat, ``Sequential
  outlier detection based on incremental decision trees,'' \emph{IEEE
  Transactions on Signal Processing}, vol.~67, no.~4, pp. 993--1005, 2018.

\bibitem{arlinghaus1994practical}
S.~Arlinghaus, \emph{Practical handbook of curve fitting}.\hskip 1em plus 0.5em
  minus 0.4em\relax CRC press, 1994.

\bibitem{buja1989linear}
A.~Buja, T.~Hastie, and R.~Tibshirani, ``Linear smoothers and additive
  models,'' \emph{The Annals of Statistics}, pp. 453--510, 1989.

\bibitem{oppenheim1997signals}
A.~V. Oppenheim, A.~S. Willsky, S.~H. Nawab, G.~M. Hern{\'a}ndez \emph{et~al.},
  \emph{Signals \& systems}.\hskip 1em plus 0.5em minus 0.4em\relax Pearson
  Educaci{\'o}n, 1997.

\bibitem{chatfield2013analysis}
C.~Chatfield, \emph{The analysis of time series: theory and practice}.\hskip
  1em plus 0.5em minus 0.4em\relax Springer, 2013.

\bibitem{wei2006time}
W.~W. Wei, ``Time series analysis,'' in \emph{The Oxford Handbook of
  Quantitative Methods in Psychology: Vol. 2}, 2006.

\bibitem{gokcesu2018adaptive}
K.~Gokcesu, M.~Ergeneci, E.~Ertan, and H.~Gokcesu, ``An adaptive algorithm for
  online interference cancellation in emg sensors,'' \emph{IEEE Sensors
  Journal}, vol.~19, no.~1, pp. 214--223, 2018.

\bibitem{hansun2013new}
S.~Hansun, ``A new approach of moving average method in time series analysis,''
  in \emph{2013 conference on new media studies (CoNMedia)}.\hskip 1em plus
  0.5em minus 0.4em\relax IEEE, 2013, pp. 1--4.

\bibitem{kaiser1977data}
J.~Kaiser and W.~Reed, ``Data smoothing using low-pass digital filters,''
  \emph{Review of Scientific Instruments}, vol.~48, no.~11, pp. 1447--1457,
  1977.

\bibitem{hatchett2010optimal}
R.~B. Hatchett, B.~W. Brorsen, and K.~B. Anderson, ``Optimal length of moving
  average to forecast futures basis,'' \emph{Journal of Agricultural and
  Resource Economics}, pp. 18--33, 2010.

\bibitem{chiarella2006dynamic}
C.~Chiarella, X.-Z. He, and C.~Hommes, ``A dynamic analysis of moving average
  rules,'' \emph{Journal of Economic Dynamics and Control}, vol.~30, no. 9-10,
  pp. 1729--1753, 2006.

\bibitem{gokcesu2021nonparametric}
K.~Gokcesu and H.~Gokcesu, ``Nonparametric extrema analysis in time series for
  envelope extraction, peak detection and clustering,'' \emph{arXiv preprint
  arXiv:2109.02082}, 2021.

\bibitem{schwert1987effects}
G.~W. Schwert, ``Effects of model specification on tests for unit roots in
  macroeconomic data,'' \emph{Journal of monetary economics}, vol.~20, no.~1,
  pp. 73--103, 1987.

\bibitem{bai2004combination}
Y.-W. Bai, W.-Y. Chu, C.-Y. Chen, Y.-T. Lee, Y.-C. Tsai, and C.-H. Tsai, ``The
  combination of kaiser window and moving average for the low-pass filtering of
  the remote ecg signals,'' in \emph{Proceedings. 17th IEEE Symposium on
  Computer-Based Medical Systems}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE,
  2004, pp. 273--278.

\bibitem{chen2003moving}
H.~Chen and S.-W. Chen, ``A moving average based filtering system with its
  application to real-time qrs detection,'' in \emph{Computers in Cardiology,
  2003}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2003, pp. 585--588.

\bibitem{justusson1981median}
B.~Justusson, ``Median filtering: Statistical properties,'' in
  \emph{Two-Dimensional Digital Signal Prcessing II}.\hskip 1em plus 0.5em
  minus 0.4em\relax Springer, 1981, pp. 161--196.

\bibitem{kalman1960new}
R.~E. Kalman and Others, ``A new approach to linear filtering and prediction
  problems,'' \emph{Journal of basic Engineering}, vol.~82, no.~1, pp. 35--45,
  1960.

\bibitem{gokcesu2021generalized}
K.~Gokcesu and H.~Gokcesu, ``Generalized huber loss for robust learning and its
  efficient minimization for a robust statistics,'' \emph{arXiv preprint
  arXiv:2108.12627}, 2021.

\bibitem{khargonekar1985non}
P.~Khargonekar and A.~Tannenbaum, ``Non-euclidian metrics and the robust
  stabilization of systems with parameter uncertainty,'' \emph{IEEE
  Transactions on Automatic Control}, vol.~30, no.~10, pp. 1005--1013, 1985.

\bibitem{gokcesu2022nonconvex}
K.~Gokcesu and H.~Gokcesu, ``Nonconvex extension of generalized huber loss for
  robust learning and pseudo-mode statistics,'' \emph{arXiv preprint
  arXiv:2202.11141}, 2022.

\bibitem{pugh1990skip}
W.~Pugh, ``Skip lists: a probabilistic alternative to balanced trees,''
  \emph{Communications of the ACM}, vol.~33, no.~6, pp. 668--676, 1990.

\bibitem{arce2005nonlinear}
G.~R. Arce, \emph{Nonlinear signal processing: a statistical approach}.\hskip
  1em plus 0.5em minus 0.4em\relax John Wiley \& Sons, 2005.

\bibitem{do1995applying}
C.~L. do~Lago, V.~F. Juliano, and C.~Kascheres, ``Applying moving median
  digital filter to mass spectrometry and potentiometric titration,''
  \emph{Analytica Chimica Acta}, vol. 310, no.~2, pp. 281--288, 1995.

\bibitem{huang1998computing}
D.~Huang and W.~T. Dunsmuir, ``Computing joint distributions of 2d moving
  median filters with applications to detection of edges,'' \emph{IEEE
  transactions on pattern analysis and machine intelligence}, vol.~20, no.~3,
  pp. 340--343, 1998.

\bibitem{ataman1981some}
E.~Ataman, V.~Aatre, and K.~Wong, ``Some statistical properties of median
  filters,'' \emph{IEEE Transactions on Acoustics, Speech, and Signal
  Processing}, vol.~29, no.~5, pp. 1073--1075, 1981.

\bibitem{weisstein2002crc}
E.~W. Weisstein, \emph{CRC concise encyclopedia of mathematics}.\hskip 1em plus
  0.5em minus 0.4em\relax Chapman and Hall/CRC, 2002.

\bibitem{rudge1982handbook}
A.~W. Rudge and K.~Milne, \emph{The handbook of antenna design}.\hskip 1em plus
  0.5em minus 0.4em\relax Iet, 1982, vol.~16.

\bibitem{stoica2005spectral}
P.~Stoica and R.~Moses, \emph{Spectral Analysis of Signals}.\hskip 1em plus
  0.5em minus 0.4em\relax Pearson Prentice Hall, 2005.

\bibitem{van1988beamforming}
B.~D. Van~Veen and K.~M. Buckley, ``Beamforming: A versatile approach to
  spatial filtering,'' \emph{IEEE assp magazine}, vol.~5, no.~2, pp. 4--24,
  1988.

\bibitem{dixon1951introduction}
W.~J. Dixon and F.~J. Massey~Jr, \emph{Introduction to statistical
  analysis.}\hskip 1em plus 0.5em minus 0.4em\relax McGraw-Hill, 1951.

\bibitem{keerthi2003asymptotic}
S.~S. Keerthi and C.-J. Lin, ``Asymptotic behaviors of support vector machines
  with gaussian kernel,'' \emph{Neural computation}, vol.~15, no.~7, pp.
  1667--1689, 2003.

\bibitem{ghosh2006introduction}
J.~K. Ghosh, M.~Delampady, and T.~Samanta, \emph{An introduction to Bayesian
  analysis: theory and methods}.\hskip 1em plus 0.5em minus 0.4em\relax
  Springer, 2006, vol. 725.

\bibitem{o1978curve}
A.~O'Hagan, ``Curve fitting and optimal design for prediction,'' \emph{Journal
  of the Royal Statistical Society: Series B (Methodological)}, vol.~40, no.~1,
  pp. 1--24, 1978.

\bibitem{wand1994kernel}
M.~P. Wand and M.~C. Jones, \emph{Kernel smoothing}.\hskip 1em plus 0.5em minus
  0.4em\relax CRC press, 1994.

\bibitem{gokcesu2022smoothing}
K.~Gokcesu and H.~Gokcesu, ``Smoothing with the best rectangle window is
  optimal for all tapered rectangle windows,'' \emph{arXiv preprint
  arXiv:2203.02997}, 2022.

\end{thebibliography}

\documentclass[journal]{IEEEtran}
\usepackage{latexsym,,amssymb,amsmath,graphicx,epsf,cite,bbm,float}
\usepackage{ifpdf}
\usepackage{epstopdf,mathtools}

\usepackage{algorithm,algorithmic}
\usepackage{amsmath,amssymb,bm}
\usepackage{amsfonts,dsfont,color,bbm,subcaption}

\newcommand{\abs}[1]{|#1|}

\usepackage{hyperref,amsthm}

\newtheorem{theorem}{Theorem}

\newtheorem{lemma}[]{Lemma}

\newtheorem{proposition}[]{Proposition}

\newtheorem{remark}[]{Remark}

\begin{document}

\title{An Auto-Regressive Formulation for Smoothing and Moving Mean with Exponentially Tapered Windows} 
\author{\IEEEauthorblockN{Kaan Gokcesu}, \IEEEauthorblockN{Hakan Gokcesu} }
\maketitle

\begin{abstract}
	We investigate an auto-regressive formulation for the problem of smoothing time-series by manipulating the inherent objective function of the traditional moving mean smoothers. Not only the auto-regressive smoothers enforce a higher degree of smoothing, they are just as efficient as the traditional moving means and can be optimized accordingly with respect to the input dataset. Interestingly, the auto-regressive models result in moving means with exponentially tapered windows.
\end{abstract}

\section{Introduction}
Data smoothing is heavily utilized in many learning problems to detect certain patters in the data or mitigate noisy/anomalous values \cite{hardle1991smoothing}. The notion of smoothing stems from the inference that the data points close to each other in their features should have close mappings (functional evaluations), i.e., behave similarly. The most prominent example is a time-series dataset where data points closer in time should have similar values, e.g, unnaturally high or low values should be suppressed to match the nominal data pattern.
Smoothing is mainly done on the datasets for easier extraction of their information and consequently ease their analyses \cite{simonoff2012smoothing}. This approach is applied in many fields including signal processing \cite{roberts1987digital,orfanidis1995introduction,fong2002monte,schaub2018flow}, data analysis \cite{kenny1998data,brandt1998data,guo2004functional,deveaux1999applied,zin2020smoothing}, machine learning \cite{gokcesu2021regret,saul2004overview,gokcesu2020recursive,yaroshinsky2001smooth,gokcesu2020generalized,servedio2003smooth,gokcesu2021optimal,hartikainen2010kalman, gokcesu2021optimally}, and anomaly detection \cite{chandola2009anomaly,gokcesu2017online,tsopelakos2019sequential, gokcesu2018sequential}.
While curve-fitting is closely related to smoothing, it has a major distinction \cite{arlinghaus1994practical}. The curve-fitting aims to construct a well-defined function using the observed data points, e.g., interpolation, which is not of interest in smoothing. 

The most traditional way of smoothing is via linear smoothers, in which each data point is smoothed with linear weighting of nearby data points \cite{buja1989linear}. In other words, given an input $\boldsymbol{y}$ and linear weights or Finite Impulse Response (FIR) filter $\boldsymbol{w}$, the smoothed $\boldsymbol{x}$ is given by their convolution \cite{oppenheim1997signals}.

The most commonly utilized linear smoother (especially for smoothing of a time-series) is the moving (rolling) mean (average) filter, where the linear weights $\boldsymbol{w}$ are all non-negative and sum to $1$ (i.e., a probability distribution) \cite{chatfield2013analysis,wei2006time,gokcesu2018adaptive,hansun2013new}. 
Hence, the weights constitute a low-pass filter to mitigate the higher frequency deviations and smooth out the original input signal \cite{kaiser1977data}.
The moving mean smoother is applied to better detect long-term behavior or trend, where the emphasize is directly related to the filter length \cite{hatchett2010optimal}.
The use of the moving mean linear smoother is commonly encountered in mathematical finance to analyze features like the trading price and volumes in the stock market \cite{chiarella2006dynamic,gokcesu2021nonparametric}. Similarly, it is also used to analyze the trends of many macroeconomic features including import/export volume, unemployment statistics and gross domestic product \cite{schwert1987effects}. It is also utilized in biomedical signal processing such as electrocardiogram \cite{bai2004combination,chen2003moving}.

Furthermore, the smoothing of an input signal is not only achieved by the moving mean filter but also other diverse filters as well, such as the rolling median filters \cite{justusson1981median}. Note that the moving mean filter becomes statistically optimal in the maximum likelihood sense when the input signal is contaminated by Gaussian noise \cite{kalman1960new}. Nonetheless, when the contamination is non-Gaussian and has heavy tails in its distributions, the performance of the moving mean degrades substantially and becomes easily influenced by undesired outliers. More robust approaches are studied to this effect \cite{gokcesu2021generalized,khargonekar1985non,gokcesu2022nonconvex}, where a prominent example is the moving median filter. Although the moving median is a nonlinear smoother and the smoothed signal is acquired by the weighted median of close groups of samples, it is efficiently computable by use of indexable skiplist \cite{pugh1990skip}. We point out that specifically for the moving median, the smoothed signal is statistically optimal in the maximum likelihood sense, when the signal contamination results from a Laplace distribution \cite{arce2005nonlinear}. The moving median filter has many applications including mass spectrometry \cite{do1995applying}, edge detection \cite{huang1998computing}. The application is especially rich in the field of image processing because of the edge-preserving capabilities \cite{ataman1981some}. 

To recap, given the observed input $\boldsymbol{y}$, the smoothed output $\boldsymbol{x}$ is generally created by use of a moving operation $\mathcal{T}(\cdot)$, i.e.,
\begin{align}
	\boldsymbol{x}=\mathcal{T}(\boldsymbol{y}),
\end{align}
where $\mathcal{T}(\cdot)$ can be either the rolling mean, the rolling median or some other function. By incorporation of weighted window function $\boldsymbol{w}$, we can generalize the operation as
\begin{align}
	\boldsymbol{x}=\mathcal{T}(\boldsymbol{y},\boldsymbol{w}),
\end{align}
where the relevant operations instead become weighted moving mean or weighted moving median, etc. In this work, we specifically focus on the moving mean smoothing operation.

We point out that in this smoothing operation, the design boils down to the weighting, which are provided by specific window functions \cite{weisstein2002crc} that are heavily utilized in many fields including antenna design \cite{rudge1982handbook}, spectral analysis \cite{stoica2005spectral} and beam-forming \cite{van1988beamforming}.
We focus on their applications in statistical analysis and smoothing \cite{dixon1951introduction}. Note that, these weighting windows are also referred as kernel functions \cite{keerthi2003asymptotic} in Bayesian analysis \cite{ghosh2006introduction} and curve fitting \cite{o1978curve}. 
Moreover, in kernel smoothing \cite{wand1994kernel}, the estimates are created from the weighted averages of nearby data points using the appropriate kernel, where the relevant weighting is a tapered window, i.e., it diminishes from the middle peak. While the traditional problem of moving mean tapering window design has been studied \cite{gokcesu2022smoothing}, we focus on its extension where the moving mean can be expressed as an auto-regressive model.


\section{Preliminaries}\label{sec:problem}

Let us have the observed samples 
\begin{align}
	\boldsymbol{y}=\{y_n\}_{n=1}^{N}.
\end{align}
We smooth $\boldsymbol{y}$ and create our smoothed signal
\begin{align}
	\boldsymbol{x}=\{x_n\}_{n=1}^{N}.
\end{align}

Traditionally, the most popular smoothing technique is the weighted moving average smoothing \cite{gokcesu2022smoothing}, where  $\boldsymbol{x}$ is created by passing $\boldsymbol{y}$ through a weighted moving average (weighted mean) filter $\boldsymbol{w}=\{w_k\}_{k=-M}^{M}$ of window size $2M+1\leq N$, where $\boldsymbol{w}$ is in a probability simplex, hence, the weights are positive, i.e.,
\begin{align}
	w_k\geq 0,&&k\in\{-M,\ldots,M\},
\end{align}
and they sum to $1$, i.e.,
\begin{align}
	\sum_{k=-M}^{M}w_k=1.
\end{align}
Let $N$ be odd, i.e., $N=2K+1$ for some natural number $K$. Then, without loss of generality, we can define the weight vector $\boldsymbol{w}$ for a window length of $N=2K+1$ with some trailing zeros, i.e.,
\begin{align}
	\tilde{w}_k=\begin{cases}
		w_k,&-M\leq k\leq M\\
		0,& \text{otherwise}
	\end{cases},
\end{align} 
for $k\in\{-K,\ldots,K\}$. Thus, without loss of generality, we can assume the window length is $N$.
The smoothed signal $\boldsymbol{x}$ is given by
\begin{align}
	x_n=\sum_{k=-K}^{K}w_ky_{n+k},\label{eq:wmean}
\end{align}  
where the mean filter is cyclic, i.e., 
\begin{align}
	y_{n'}=y_{n'+N}, &&\forall n'.\label{eq:ycyclic}
\end{align}
Note that the output $\boldsymbol{x}$ of this weighted moving average is a global minimizer of the following optimization problem (weighted cumulative cross error):
\begin{align}
	\min_{\boldsymbol{x}\in\Re^N}\sum_{n=1}^{N}\sum_{k=-K}^{K}w_k(y_{n+k}-x_n)^2,\label{eq:probx}
\end{align} 
since it is convex in $\boldsymbol{x}$ and the gradient is zero at its minimizer. Thus, given a weight vector $\boldsymbol{w}$ and the input $\boldsymbol{y}$, the solution of \eqref{eq:probx} is given by the $\boldsymbol{x}$ in \eqref{eq:wmean}. 

In this paper, instead of the objective in \eqref{eq:probx}, we study the following revised objective function to better enforce the smoothness in our output $\boldsymbol{x}$:

\begin{align}
F(\boldsymbol{x},\boldsymbol{y},\boldsymbol{w})=\sum_{n=1}^{N}\sum_{k=-K}^{K}w_k(z_{n,k}-x_n)^2,\label{eq:probxS}
\end{align} 
where 
\begin{align}
	z_{n,k}=\begin{cases}
		x_{n+k},& k\neq 0\\
		y_n, & k=0 
	\end{cases}.
\end{align}

We can extend the objective function to the following generalized form:

\begin{align}
	G(\boldsymbol{x,y,\alpha,\beta})=\sum_{n,k}\alpha_k(y_{n+k}-x_n)^2+\sum_{n,k}\beta_k(x_{n+k}-x_n)^2,\label{objG}
\end{align} 
for some nonnegative weights $\alpha_k$ and $\beta_k$, i.e.,
\begin{align}
	\alpha_k\geq& 0, &&\forall k,
	\\\beta_k\geq& 0, &&\forall k.
\end{align}

\begin{proposition}\label{thm:eqForm}
	The objective function in \eqref{objG} is equivalent to
	\begin{align*}
		G(\boldsymbol{x,y,\alpha,\beta})=H_0(\boldsymbol{y,\alpha})+H(\boldsymbol{x,\bar{y},\beta},A),
	\end{align*}
	where
	\begin{align*}
		H_0(\boldsymbol{y,\alpha})=&\sum_{n,k}\alpha_k(y_{n+k}-\bar{y}_n)^2,
		\\H(\boldsymbol{x,\bar{y},\beta},A)=&\sum_{n}A(\bar{y}_{n}-x_n)^2+\sum_{n,k}\beta_k(x_{n+k}-x_n)^2,
		\\\bar{y}_n=&\frac{\sum_{k}\alpha_ky_{n+k}}{A},
		\\A=&\sum_{k}\alpha_k.
	\end{align*}

	\begin{proof}
		The proof comes from the principle of orthogonality. We have
		\begin{align}
			\sum_{k}\alpha_k(y_{n+k}-x_n)^2=&\sum_{k}\alpha_k(y_{n+k}-\bar{y}_n+\bar{y}_n-x_n)^2,
			\\=&\sum_{k}\alpha_k\left[(y_{n+k}-\bar{y}_n)^2+(\bar{y}_n-x_n)^2\right]\nonumber
			\\&+\sum_{k}2\alpha_k(y_{n+k}-\bar{y}_n)(\bar{y}_n-x_n),
			\\=&\sum_{k}\alpha_k\left[(y_{n+k}-\bar{y}_n)^2+(\bar{y}_n-x_n)^2\right].
		\end{align}
		which concludes the proof.
	\end{proof}
\end{proposition}

\begin{remark}\label{thm:eqForm2}
	The objective function $H(\boldsymbol{x,\bar{y},\beta},A)$ in \autoref{thm:eqForm} is equivalent to the objective function $F(\boldsymbol{x,y,w})$ in \eqref{eq:probxS}, i.e.,
	\begin{align*}
		H(\boldsymbol{x,\bar{y},\beta},A)\leftrightarrow F(\boldsymbol{x,y,w})
	\end{align*}
	when
	\begin{align*}
		\bar{y}&\leftrightarrow y
		\\w_0&\leftrightarrow \frac{A}{A+B}
		\\w_k&\leftrightarrow\frac{\beta_k}{A+B}, &&k\neq0
	\end{align*}
\end{remark}

Thus, without loss of generality, we have the objective

\begin{align}
	\min_{\boldsymbol{x}\in\Re^N}\left[\sum_{n}w_0(\bar{y}_{n}-x_n)^2+\sum_{n}\sum_{k\neq 0}w_k(x_{n+k}-x_n)^2,\right],
\end{align} 
i.e., we want to minimize $F(\boldsymbol{x,\bar{y},w})$.

\section{Extracting the Smoothed Signal}\label{sec:fft}

In this section, we extract the smoothed signal $\boldsymbol{x}$.

\begin{lemma}\label{thm:convex}
	The objective function $F(\boldsymbol{x,\bar{y},w})$ in \eqref{eq:probxS} is convex in its argument $\boldsymbol{x}$.
	
	\begin{proof}
		Let $x_n=\lambda x_{n}^{(1)}+(1-\lambda)x_{n}^{(2)}$.
		We have
		\begin{align}
			\left(\bar{y}_{n}-x_n\right)^2=&\left(\bar{y}_{n}-\lambda x_{n}^{(1)}-(1-\lambda)x_{n}^{(2)}\right)^2,
			\\=&\left(\lambda\left(\bar{y}_{n}-x_{n}^{(1)}\right)+(1-\lambda)\left(\bar{y}_{n}-x_{n}^{(2)}\right)\right)^2,\nonumber
			\\\leq&\lambda\left(\bar{y}_{n}-x_{n}^{(1)}\right)^2+(1-\lambda)\left(\bar{y}_{n}-x_{n}^{(2)}\right)^2,
		\end{align}
		since the square function is convex. Similarly,
		\begin{align*}
			(x_{n+k}-x_n)^2=&\left[\lambda\left[ x_{n+k}^{(1)}-x_{n}^{(1)}\right]+(1-\lambda)\left[x_{n+k}^{(2)}-x_{n}^{(2)}\right]\right]^2,
			\\\leq&\lambda\left( x_{n+k}^{(1)}-x_{n}^{(1)}\right)^2+(1-\lambda)\left(x_{n+k}^{(2)}-x_{n}^{(2)}\right)^2.
		\end{align*}
		Thus, we have
		\begin{align}
			w_0&(\bar{y}_{n}-x_n)^2+\sum_{k\neq 0}w_k(x_{n+k}-x_n)^2
			\\\leq&\lambda\left[w_0\left(\bar{y}_{n}-x_n^{(1)}\right)^2+\sum_{k\neq 0}w_k\left(x_{n+k}^{(1)}-x_n^{(1)}\right)^2\right]\nonumber
			\\&+(1-\lambda)\left[w_0\left(\bar{y}_{n}-x_n^{(2)}\right)^2+\sum_{k\neq 0}w_k\left(x_{n+k}^{(2)}-x_n^{(2)}\right)^2\right].\nonumber
		\end{align}
		Hence, 
		\begin{align}
			F(\boldsymbol{x,\bar{y},w})\leq\lambda F(\boldsymbol{x^{(1)},\bar{y},w})+(1-\lambda)F(\boldsymbol{x^{(2)},\bar{y},w}),\nonumber
		\end{align}
		which concludes the proof.
	\end{proof}
\end{lemma}

We observe that because of the inherent symmetry in the objective function $F(\boldsymbol{x,\bar{y},w})$ due to the summation over $n$, we can manipulate $\boldsymbol{w}$ such that it is symmetric, i.e.,
\begin{align}
	w_k=w_{-k}, \forall k.
\end{align}

\begin{lemma}\label{thm:conv}
	The minimizer $\boldsymbol{x^*}$ of $F(\boldsymbol{x,\bar{y},w})$ is such that the circular convolution of $\boldsymbol{x}=\{x_n\}_{n=1}^N$ and $\boldsymbol{v}=\{v_k\}_{k=-K}^K$ gives $\boldsymbol{\bar{y}}=\{\bar{y}_n\}_{n=1}^N$, i.e., for positive $w_0$,
	\begin{align*}
		\boldsymbol{v}\circledast\boldsymbol{x}=\boldsymbol{\bar{y}},&&v_k=\begin{cases}
			\frac{2}{w_0}-1,& k=0\\
			-\frac{2w_k}{w_0},& k\neq0
		\end{cases}.
	\end{align*}
	\begin{proof}
		Since the objective function $F(\boldsymbol{x,\bar{y},w})$ is convex in $\boldsymbol{x}$ from \autoref{thm:convex}, its derivative at a minimizer $\boldsymbol{x_*}$ should be zero. Thus, since $w_k$ is symmetric and $\sum_{k}w_k=1$, we have
		\begin{align}
			(2-w_0)x_n=&w_0\bar{y}+\sum_{k\neq 0}w_kx_{n+k}+\sum_{k\neq 0}w_kx_{n-k},
			\\=&w_0\bar{y}+2\sum_{k\neq 0}w_kx_{n+k},
		\end{align}
		If $w_0=0$, any constant $\boldsymbol{x}$ is a solution. If $w_0>0$, we have
		\begin{align}
			\sum_{k}v_kx_{n+k}=\bar{y},
		\end{align}
		which concludes the proof.
	\end{proof}
\end{lemma}

To extract $\boldsymbol{x}$, we need to deconvolve $\boldsymbol{\bar{y}}$ with $\boldsymbol{v}$.

\begin{lemma}\label{thm:inv}
	The convolution by $\boldsymbol{v}$ is reversible, i.e., there exists a unique deconvolution.
	\begin{proof}
		Let $\mathcal{F}(\cdot)$ be the Fourier (DFT) operator. Because of the circular convolution, we have
		\begin{align}
			\mathcal{F}(\boldsymbol{v})\otimes\mathcal{F}(\boldsymbol{x})=\mathcal{F}(\boldsymbol{\bar{y}}),
		\end{align}
		where $\otimes$ denotes the Hadamard multiplication, i.e., element-wise multiplication.
		Let $\boldsymbol{V}$ be the DFT of $\boldsymbol{v}$, i.e., $\mathcal{F}(\boldsymbol{v})=\boldsymbol{V}$. Thus, the DFT of $\boldsymbol{v}$ is given by
		\begin{align}
			\boldsymbol{V}_n=&\sum_{k}v_ke^{-j\frac{2\pi kn}{N}},
			\\=&v_0+\sum_{k\neq 0}v_ke^{-j\frac{2\pi kn}{N}}
			\\=&v_0+\sum_{k=1}^Kv_k(e^{-j\frac{2\pi kn}{N}}+e^{j\frac{2\pi kn}{N}}),
		\end{align}
		because of the symmetry of $v_k$. Hence,
		\begin{align}
			\boldsymbol{V}_n=&v_0+\sum_{k=1}^K2v_k\cos\left(\frac{2\pi kn}{N}\right),
			\\=&1-\sum_{k=1}^{K}2v_k+\sum_{k=1}^K2v_k\cos\left(\frac{2\pi kn}{N}\right),
		\end{align}
		since $\sum_{k}v_k=1$. Thus,
		\begin{align}
			\boldsymbol{V}_n
			=&1-\sum_{k=1}^K2v_k\left(1-\cos\left(\frac{2\pi kn}{N}\right)\right),
			\\=&1+\sum_{k=1}^K 4\frac{w_k}{w_0}\left(1-\cos\left(\frac{2\pi kn}{N}\right)\right),
			\\\geq&1
		\end{align}
		since $w_k\geq 0$, $w_0>0$ and $\cos(\theta)\leq 1, \forall\theta$. Since all DFT points of $\boldsymbol{V}$ are nonzero, its element-wise multiplication is invertible, which concludes the proof.
	\end{proof}
\end{lemma}

\begin{theorem}
	Given the formulation in \eqref{objG} with the weights $\boldsymbol{\alpha}=\{\alpha_k\}_{k=-K}^K$ ($A=\sum_{k}\alpha_k$), $\boldsymbol{\beta}=\{\beta_k\}_{k=-K}^K$ ($B=\sum_{k}\beta_k$) and input $\boldsymbol{y}=\{y_n\}_{n=1}^K$; the minimizer $\boldsymbol{x_*}$ is
	\begin{align*}
		\boldsymbol{x_*}=\mathcal{F}^{-1}\left(\mathcal{F}\left(\boldsymbol{y}\right)\otimes\mathcal{F}(\boldsymbol{p})\oslash\mathcal{F}(\boldsymbol{v})\right),
	\end{align*}
where $\oslash$ denotes the Hadamard division, i.e., element-wise division, and
\begin{align*}
	p_k=\frac{1}{A}\alpha_k&& v_k=\begin{cases}
		\frac{A+2B}{A},& k=0\\
		-\frac{2\beta_k}{A},& k\neq0
	\end{cases}.
\end{align*}
	\begin{proof}
		From \autoref{thm:eqForm}, we have
		\begin{align}
			\mathcal{F}(\boldsymbol{\bar{y}})=\mathcal{F}(\boldsymbol{y})\otimes\mathcal{F}\left(\boldsymbol{p}\right).
		\end{align}
		 From \autoref{thm:conv} and \autoref{thm:inv}, we have		
		\begin{align}
			\mathcal{F}(\boldsymbol{x_*})=\mathcal{F}(\boldsymbol{\bar{y}})\oslash\mathcal{F}(\boldsymbol{v}),
		\end{align} 
		which concludes the proof.
	\end{proof}
\end{theorem}

\section{Weight Design}

Given the general objective in \eqref{objG}, after extracting the smoothed signal $\boldsymbol{x}$ as in \autoref{sec:fft}, the question remains as to which weights are best suited to be chosen for optimal smoothing. Let 
\begin{align}
	J(\boldsymbol{y,\theta})=\min_{\boldsymbol{x}}G(\boldsymbol{x,y,\alpha,\beta}),
\end{align}
where $\boldsymbol{\theta}=(\boldsymbol{\alpha,\beta})$ cumulatively represents the weights. 

To model the optimal smoothing we can minimize the objective function $J(\boldsymbol{y,\theta})$. Since the weights $\boldsymbol{\alpha,\beta}$ are nonnegative, i.e., $\boldsymbol{\alpha}\geq 0$ and $\boldsymbol{\beta}\geq 0$, the minimizer weights $\boldsymbol{\theta_*}$ would be all-zero because the objective is always nonnegative. To circumvent this, we need to restrict the weights.
\begin{remark}
	We use the most popular regularization, which is to limit the sum of weights to $1$, i.e., $\abs{\boldsymbol{\theta}}_1=1$ and represents a probability distribution. Hence, $A+B=1$.  
\end{remark}

\begin{lemma}
	$J(\boldsymbol{y,\theta})$ is concave in $\boldsymbol{\theta}$.
	\begin{proof}
		$G(\boldsymbol{x,y,\alpha,\beta})$ is linear in the weights $\boldsymbol{\theta}=(\boldsymbol{\alpha,\beta})$, i.e., 
		\begin{align}
			G(\boldsymbol{x,y},\sigma\boldsymbol{\alpha},\sigma\boldsymbol{\beta})=\sigma G(\boldsymbol{x,y},\boldsymbol{\alpha},\boldsymbol{\beta}),
		\end{align}
		for $\forall \sigma\in\Re$, and $J(\boldsymbol{y,\theta})$ is its minimization with respect to $\boldsymbol{x}$; which concludes the proof from \cite{gokcesu2022smoothing}.
	\end{proof}
\end{lemma}

Let $\tilde{\alpha}_n=\alpha_n/A$ and $\tilde{\beta}_n=\beta_n/B$. We can see that each of $\boldsymbol{\theta_\alpha}=(\boldsymbol{\tilde{\alpha},0})$ and $\boldsymbol{\theta_\beta}=(\boldsymbol{0,\tilde{\beta}})$ are probability distributions themselves, where the original weighting distribution is given by their convex combination $\boldsymbol{\theta}=A\boldsymbol{\theta_\alpha}+B\boldsymbol{\theta_\beta}$.
An optimal set of weights for a concave function in a convex polytope is one of its vertices \cite{gokcesu2022smoothing}. Hence, the unrestrained minimization results in either all zeros for $\boldsymbol{\alpha}$ or all zeros for $\boldsymbol{\beta}$. If $\boldsymbol{\beta}$ is all-zeros, the smoothed signal will be equal to the weighted mean of $\boldsymbol{y}$ with $\boldsymbol{\alpha}$. If $\boldsymbol{\alpha}$ is all-zeros, the smoothed signal will be an all-constant signal, which will result in the minimum error (trivially zero). Hence, increasing $B$ directly decreases the error.

To circumvent these trivialities, we need to restrain $A$ and $B$. Given fixed $A$ and $B$ (where $A+B=1$), we have
\begin{align}
	p_k=&\frac{\alpha_k}{A}, \\q_k=&\frac{\beta_k}{B},
	\\v_k=&\begin{cases}
		\frac{2}{A}-1,& k=0\\
		-\frac{2B}{A}q_k,& k\neq0
	\end{cases}
\end{align}
Unrestrained optimization of $\boldsymbol{p}$ and $\boldsymbol{q}$ will again result in a trivial solution. To this end, we restrict them to symmetric tapering window functions such that
\begin{align}
	p_{-1}=&p_0=p_1
	\\p_k=&p_{-k}
	\\p_{k_1}\geq& p_{k_2}, &&\abs{k_1}<\abs{k_2}
\end{align}
Similarly, for $q_k$,
\begin{align}
	q_k=&q_{-k}
	\\q_{k_1}\geq& q_{k_2}, &&\abs{k_1}<\abs{k_2}
\end{align}

\begin{remark}
	$A=1/3$ and $B=2/3$ is a natural choice to guarantee tapering behavior for $w_k$ as in \autoref{thm:eqForm2} while making $A$ as small as possible. 
\end{remark}
We observe that both $p_k$ and $q_k$ can be written in terms of a convex mixture of bandlimited uniform distributions because of the tapering. The convex mixture of uniform distributions create a convex polytope. Because of the concavity of the objective, one of the vertices is the solution \cite{gokcesu2022smoothing}.
\begin{remark}
	Given a window length limitation $O(L)$ for both $p_k$ and $q_k$; there exist $O(L^2)$ vertices in this polytope. We can try all of these vertices to find the optimizer, which will take $O(L^2N\log N)$ time. If $L$ is polynomially dependent on $\log N$, then the time complexity will be quasilinear, i.e., $\tilde{O}(N)$.
\end{remark}
\begin{remark}
	For increased efficiency, we may also limit the distributions to equal to each other, which would result in a polytope with $O(L)$ vertices. Consequently, the time complexity will be $O(LN\log N)$ and similarly $\tilde{O}(N)$ for poly-logarithmic $L$. 
\end{remark}
\begin{remark}
	If we do not want to limit the choice of $A$ and $B$, we may do a heuristic optimization in a cascade manner. We can first optimize $\boldsymbol{p}$ using $H_0(\boldsymbol{y,\alpha})$ in \autoref{thm:eqForm}, which will give us $\boldsymbol{\bar{y}}$.
	From $\boldsymbol{\bar{y}}$, we can optimize the weighting $\boldsymbol{q}$ using $H(\boldsymbol{x,\bar{y},\beta},A)$. However, only the tapering limitation will trivially result in a wide-bandwidth weighting. To this end, we can smooth with the narrowest bandwidth AR filter (length $3$), i.e., $w_{-1}=w_0=w_1=1/3$. The whole process to find $\boldsymbol{x}$ will take $O(N\log N)$ linearithmic time \cite{gokcesu2022smoothing}.
\end{remark}

\begin{remark}
	In the cascade, the first filter $\boldsymbol{p}$ is a weighted mean filter. While the second AR filter, the deconvolution by $\boldsymbol{v}$, has a curious window shape. While it is also a unimodal window function, it has exponential tapering at its asymptotes. Let us define the deconvolution with $\boldsymbol{u}$, whose convolution with $\boldsymbol{v}$ will be the dirac-delta. Hence, the elements of $\boldsymbol{u}$ will have a recurrence relation (linear difference equation) with the coefficients $\boldsymbol{v}$. Consider the characteristic polynomial implied by $\boldsymbol{v}$, which is
	\begin{align}
		\Lambda(r)=\sum_{k=-K}^{K}v_kr^k.
	\end{align}
	Since $v_k$ is symmetric, we have
	\begin{align}
		\Lambda(r)= v_0+\sum_{k=1}^{K}v_k(r^k+r^{-k}).
	\end{align}
	From \autoref{thm:inv}, we know that this polynomial does not have complex roots on the unit circle. Moreover, we have
	\begin{align}
		\lim_{r\rightarrow0^{+}}\Lambda(r)=-\infty,&& 	\Lambda(1)=v_0+\sum_{k=1}^{K}2v_k=1>0,
	\end{align}
	since $v_k\leq 0$ for $k\neq 0$. Furthermore, since
	\begin{align}
		\Lambda'(r)=\sum_{k=1}^{K}kv_kr^{-1}(r^{k}-r^{-k})>0, &&r\in(0,1),
	\end{align} 
	we have a unique root in $(0,1)$. Thus, the deconvolution window will have near exponentially decreasing tails. 
\end{remark}


\bibliographystyle{IEEEtran}
\bibliography{double_bib}

\end{document}